# Paul Wild    Lutz Schröder    Karla Messing    Barbara König    Jonas Forster

### v1

@@ Abstract
The classical Kantorovich-Rubinstein duality guarantees coincidence between metrics on the space of probability distributions defined on the one hand via transport plans (couplings) and on the other hand via price functions. Both constructions have been lifted to the level of generality of set functors, with the coupling-based construction referred to as the Wasserstein lifting, and the price-function-based construction as the Kantorovich lifting, both based on a choice of quantitative modalities for the given functor. It is known that every Wasserstein lifting can be expressed as a Kantorovich lifting; however, the latter in general needs to use additional modalities. We give an example showing that this cannot be avoided in general. We refer to cases in which the same modalities can be used as satisfying the generalized Kantorovich-Rubinstein duality. We establish the generalized Kantorovich-Rubinstein duality in this sense for two important cases: The Lévy-Prokhorov distance on distributions, which finds wide-spread applications in machine learning due to its favourable stability properties, and the standard metric on convex sets of distributions that arises by combining the Hausdorff and Kantorovich-Wasserstein distances.
@@ 1 Introduction
Measuring behavioural distances between probabilistic systems requires notions of distance between probability distributions (e.g. [29]). One well-established metric on the set of distributions over a metric space is variously termed the Kantorovich, Wasserstein, or Hutchinson metric. It can be calculated either by minimizing over the expected value of transport plans between, or couplings of, the given distributions, or by maximizing over the difference of expectations taken over all non-expansive price functions. The coincidence of these two values is the classical Kantorovich-Rubinstein duality [30, Theorem 5.10]. Intuitively speaking, a transport plan or coupling is a way to transform one distribution into another by shifting around weight, and its cost, to be minimized, is determined by how much weight is shifted over which distances. On the other hand, a price function determines a price for some commodity at given points; non-expansiveness of the price function means that no profit can be made from full-cost transport. The difference between the expected values of a price function under the given distributions is the profit to be made by having the commodity transported, and hence the amount one can offer to a logistics provider when outsourcing the transport.
@@
Both the coupling-based definition and the price-function-based definition have been generalized categorically to construct liftings of set functors to the category of (pseudo-)metric spaces [4] and quantitative lax extensions of set functors [32], where the latter are distinguished by applying to unrestricted quantitative relations instead of only to (pseudo-)metrics. Metric functor liftings and lax extensions in particular serve to give a general treatment of behavioural distances on quantitative systems such as probabilistic, weighted, or metric [2] transition systems in the framework of universal coalgebra [25]. In this framework, functors serve as parameters determining a type of systems as their coalgebras; for instance, coalgebras for the distribution functor are Markov chains. In the generalized setting, the coupling-based construction is typically referred as the Wasserstein lifting or extension, and the price-function-based one as the Kantorovich lifting or extension. Both constructions are parametrized over a choice of quantitative modalities; the classical case involves, on both sides, only one modality, the expectation modality.
@@
The interest in having a Kantorovich presentation of a given functor lifting or lax extension lies inter alia in the fact that one obtains a quantitative Hennessy-Milner property for the quantitative modal logic generated by the respective modalities [20, 32]. This property states coincidence of the behavioural distance induced by the given lifting or extension with the logical distance induced by the respective quantitative modal logic; thus, high distance between two states can always be certified by means of a modal formula, a principle that has received increased recent attention (e.g. [24, 28]). It has been shown that every metric functor lifting that preserves isometries [13] and every quantitative lax extension [32] is Kantorovich, i.e. can be presented via the generalized Kantorovich construction, using however a rather large (in particular typically infinite) set of modalities called the Moss modalities.
@@
An important open point that remains is thus the question of what we term generalized Kantorovich-Rubinstein duality: In which cases does a Wasserstein distance given by a choice of modalities for a functor coincide with the Kantorovich distance for the same modalities? Known positive examples include, as mentioned, the classical Kantorovich-Wasserstein distance on distributions, but also the Hausdorff distance on the powerset. We begin our analysis by giving the first example of a natural Wasserstein metric for which generalized Kantorovich-Rubinstein duality in this sense fails, namely the standard $p$-Wasserstein metric for $p>1$ (which minimizes $p$-th roots of the expectation of $p$-th powers of couplings). As our main contribution, we then provide two new positive examples, namely the Lévy-Prokhorov distance on probability distributions [23] and the standard distance on the convex powerset, whose elements are convex sets of probability distributions. The Lévy-Prokhorov distance has seen a recent rise in popularity due to its favourable robustness properties that make it suitable for tasks in machine learning (such as conformal prediction [3] and corruption resistance [5]); moreover, it has been shown in recent work to induce precisely the behavioural distance defined by $\varepsilon$-bisimilarity [10]. In this case, the relevant modality is precisely the generally modality used in work on fuzzy description logics as an alternative formal correspondent of the natural-language term ‘probably’ [27]. The convex powerset plays a central role in the distribution semantics of Markov decision processes (or probabilistic automata) [7]. Its standard metric is just the composition of the Hausdorff and Kantorovich-Wasserstein metrics, and as such is again given via the Wasserstein construction for a modality composed from the expectation modality and the standard fuzzy diamond modality [32]. We show that the generalized Kantorovich-Rubinstein duality holds w.r.t. this modality. Beyond the mentioned benefits regarding characteristic quantitative modal logics, we demonstrate in this case that the new Kantorovich description also allows for more efficient computation of distances.
@@
Our Kantorovich presentation of convex powerset complements earlier results on a Wasserstein characterization [32] and a quantitative-algebraic presentation [21]. There has been recent work on what are termed correspondences between the generalized Kantorovich and Wasserstein constructions where a single modality is assumed for the Wasserstein construction, while an associated set of modalities is considered for the Kantorovich construction [16]. Correspondences in this sense thus lie between dualities as considered here, where we insist on the same modalities being used on both sides, and general theorems on Kantorovich presentations of lax extensions [32] and classes of functors [13] that use very large sets of liftings. Results are obtained for functors constructed from the main known instances (distributions with the standard Kantorovich-Wasserstein metric, powerset) by applying coproduct and product. The Kantorovich construction is sometimes referred to as the codensity construction [18, 19, 16], and as such has been used for logical characterizations of behavioural distances as mentioned above [19] but also for game characterizations [18]. The problem of generalized Kantorovich-Rubinstein duality has been stated already in work introducing coalgebraic behavioural distances [4], and a simple counterexample has been given; the counterexample we give here is distinguished by involving a quantitative modality that satisfies an analogue of two-valued separation [22].
@@ 2 Preliminaries
A function between the underlying sets $f\colon X\to Y$ of two pseudometric spaces $(X,d_{X})$, $(Y,d_{Y})$ is nonexpansive if distances are not increased by $f$, explicitly if for all $x,y\in X$ it holds that $d_{Y}(f(x),f(y))\leq d_{X}(x,y)$. Pseudometric spaces and nonexpansive functions between them form a category $\mathbf{PMet}$. The full subcategories of $\mathbf{PMet}$ spanned by metric spaces is denoted by $\mathbf{Met}$.
@@
Our main results derive some of their interest from their relevance to behavioural distances in coalgebras. Generally, the framework of universal coalgebra [25] is based on abstracting state-based systems as $F$-coalgebras for s functor $F\colon\mathbf{C}\to\mathbf{C}$ on a category $\mathbf{C}$, with $F$ determining the type of the system. Specifically, an $F$-coalgebra is a pair $(C,\gamma)$ consisting of a $\mathbf{C}$ object $C$, thought of as an object of states, and a morphism $\gamma\colon C\to FC$ determining transitions from states to structured collections of successor states, with the structure determined by $F$. A homomorphism between $F$-coalgebras $(C,\gamma)$ and $(D,\delta)$ is a $\mathbf{C}$-morphism $h\colon C\to D$ such that $\delta\circ h=Fh\circ\gamma$.
We list some common functors that occur in the description of $F$ in the definition above.
@@
    $\mathcal{D}X=\{\mu\colon X\to[0,1]\mid\mu(x)=0\text{ for almost all }x\in X\text{ and }\sum_{x\in X}\mu(x)=1\}$
@@
    The coalgebras of $\mathcal{D}$ are precisely Markov chains.
@@ 3 Dual Characterizations of Metrics
    The Kantorovich-Wasserstein lifting $\overline{\mathcal{D}}\colon\mathbf{PMet}\to\mathbf{PMet}$ of $\mathcal{D}$ equips the set $\mathcal{D}X$ with the Kantorovich-Wasserstein distance $\delta^{\mathsf{K}}(d_{X})$, which is defined as
    $\delta^{\mathsf{K}}(d_{X})(\mu,\nu):=\sup\{\mathbb{E}_{\nu}({f})-\mathbb{E}_{\mu}({f})\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexpansive}\},$
@@
    Let $\lambda$ be a predicate lifting for $F$. The Kantorovich lifting of $F$ to the category of pseudometric spaces sends a metric $d_{X}$ to $K_{\lambda}(d_{X})\colon FX\times FX\to[0,1]$ defined by
@@
    Let $s,t\in FX$. The set of couplings $\Gamma(s,t)$ is defined as
    $\Gamma(s,t):=\{c\in F(X\times X)\mid F\pi_{1}(c)=s\text{ and }F\pi_{2}(c)=t\}$
@@
    Let $\lambda$ be a well-behaved predicate lifting for $F$ and assume that $F$ preserves weak pullbacks. The Wasserstein lifting of $F$ to the category of pseudometric spaces is defined as the lifting that equips $FX$ with $W_{\lambda}(d_{X})$ where
@@
While the Kantorovich lifting assumes no conditions on supplied structures, the Wasserstein lifting is significantly more particular, requiring both pullback preservation of the underlying functor and that the predicate lifting be well-behaved, to ensure that $W_{\lambda}(d_{X})$ is a pseudometric whenever $d_{X}$ is [4, 6].
@@
    Let $F=\mathcal{D}$ and let $\lambda$ be the predicate lifting calculating expected values: For $\mu\in\mathcal{D}X$ and $f\in[0,1]^{X}$ we have $\lambda_{X}(f)(\mu)=\mathbb{E}_{\mu}({f})$. Then $\delta^{\mathsf{K}}=K_{\lambda}=W_{\lambda}$.
If, like in the two examples above, the categorical Kantorovich and Wasserstein constructions $K_{\lambda}$ and $W_{\lambda}$ coincide, we say that generalized Kantorovich-Rubinstein duality holds. This name is motivated by the particular case of Example˜3.2, the classical Kantorovich-Rubinstein duality that dates back to the beginnings of transportation theory [17]. It is important to note that generalized duality in this sense may fail, as demonstrated by the case of $p$-Wasserstein distance that we discuss later. It is a general fact [4, Theorem 5.27] that $K_{\lambda}\leq W_{\lambda}$, so duality hinges on the inequality $W_{\lambda}\leq K_{\lambda}$.
@@
The Wasserstein presentation is closely related to optimal transportation theory: In the case of the Kantorovich-Wasserstein lifting, one may view probability distributions $\mu,\nu$ as producers and consumers of a resource respectively (with the amount of the resource produced/consumed being fixed to $1$). Now one wants to transport the produced resources to the consumers in the economically most efficient possible way, minimizing the average distance each unit of resource needs to travel. Couplings $c$ of $\mu$ and $\nu$ can then be seen as transport plans, with $\mathbb{E}_{c}({d_{X}})$ giving the total cost of carrying out the plan. Then the Kantorovich-Wasserstein distance gives us by definition the minimal possible cost.
In the distribution case (Example˜3.2), the Kantorovich presentation can be explained via the analogy of “outsourcing the transport” by defining a function $f$ that assigns a price to each $x\in X$. This function must satisfy the requirement that it is non-expansive, i.e., that the difference of prices assigned to $x,y$ is always at most $d_{X}(x,y)$ (meaning that no extra profit can be made from such a transport). The overall profit under such a price function is then the income obtained from the consumers ($\mathbb{E}_{\nu}({f})$) minus the cost paid to the producers ($\mathbb{E}_{\mu}({f})$). Taking the maximum over all such $f$ gives us the value of the Kantorovich lifting.
@@
The Kantorovich lifting is closely related to characteristic multi-valued modal logics: These are logics in which formulae $\varphi$ receive semantics in coalgebras $(X,\gamma)$, inducing an interpretation function $\llbracket\varphi\rrbracket_{\gamma}\colon X\to[0,1]$. The logical distance of two states is then the supremum of all distances witnessed by such formulas.
The semantics of a modal operator $L$ in these types of logics is usually given as a predicate lifting $\lambda\colon[0,1]^{-}\Rightarrow[0,1]^{F-}$, with $\llbracket L\varphi\rrbracket_{\gamma}$ being inductively defined as $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$. Then expressivity of the logic (the fact that the behavioural distance can be witnessed by formulae of the logic arbitrarily closely) can be shown by exploiting the Kantorovich-presentation of the lifting. In fact, the interpretation $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$ can be viewed as emulating one step of the functional $\Phi_{\gamma}$, under the condition that the interpretations of formulae $\llbracket\varphi\rrbracket_{\gamma}$ are able to approximate any nonexpansive function $f\colon X\to[0,1]$ arbitrarily closely [12].
Hence, the duality can be used very fruitfully: One can use the Wasserstein view to compute (an under-approximation of) the behavioural distance and switch to Kantorovich to determine the distinguishing formula witnessing this distance.
@@
There also exists a parametrized version of the Kantorovich-Wasserstein distance, called the $p$-Wasserstein distance for some real parameter $p\geq 1$}. In categorical terms, it is given by the predicate lifting $\lambda_{p}(f)(\mu)=(\mathbb{E}_{\mu}({f^{p}}))^{\frac{1}{p}}$ (where $f^{p}$ takes the $p$-th power pointwise), that is:
@@
For $p=1$ this is just the usual Kantorovich-Wasserstein distance and duality holds. If $p>1$, however, then duality may fail, and the corresponding Kantorovich construction $K_{\lambda_{p}}$ may be strictly below $W_{\lambda_{p}}$:
@@
$W_{\lambda_{2}}(d)(\mu,\nu)=(\delta^{\mathsf{K}}(d)(\mu,\nu))^{\frac{1}{2}}=(\tfrac{1}{3})^{\frac{1}{2}}=\tfrac{1}{\sqrt{3}},$
@@
$g(0)+g(1)=(\tfrac{1}{3}\cdot g(0)+\tfrac{2}{3}\cdot g(1))+(\tfrac{2}{3}\cdot g(0)+\tfrac{1}{3}\cdot g(1))\\ \leq g(\tfrac{1}{3}\cdot 0+\tfrac{2}{3}\cdot 1)+g(\tfrac{2}{3}\cdot 0+\tfrac{1}{3}\cdot 1)=g(\tfrac{1}{3})+g(\tfrac{2}{3})$ (1)
@@
Dividing by $g(\tfrac{2}{3})+g(\tfrac{1}{3})$ on both sides results in the claimed inequality. ∎

### v2

@@ Abstract
The classical Kantorovich-Rubinstein duality guarantees coincidence between metrics on the space of probability distributions defined on the one hand via transport plans (couplings) and on the other hand via price functions. Both constructions have been lifted to the level of generality of set functors, with the construction based on couplings referred to as the Wasserstein or simply the coupling-based lifting, and the price-function-based construction as the Kantorovich or codensity lifting, both based on a choice of quantitative modalities for the given functor. It is known that every coupling-based lifting can be expressed as a price-function-based lifting; however, the latter in general needs to use additional modalities. We give an example showing that this cannot be avoided in general. We refer to cases in which the same modalities can be used as satisfying the generalized Kantorovich-Rubinstein duality. We establish the generalized Kantorovich-Rubinstein duality in this sense for two important cases: The Lévy-Prokhorov distance on distributions, which finds wide-spread applications in machine learning due to its favourable stability properties, and the standard metric on convex sets of distributions that arises by combining the Hausdorff and Kantorovich-Rubinstein distances.
@@ 1 Introduction
Measuring behavioural distances between probabilistic systems requires notions of distance between probability distributions (e.g. [36]). One well-established metric on the set of distributions over a metric space is variously termed the Kantorovich-Rubinstein, Wasserstein, or Hutchinson metric. It can be calculated either by minimizing over the expected value of transport plans between, or couplings of, the given distributions, or by maximizing over the difference of expectations taken over all nonexpansive price functions. The coincidence of these two values is the classical Kantorovich-Rubinstein duality [37, Theorem 5.10]. Intuitively speaking, a transport plan or coupling is a way to transform one distribution into another by shifting around weight, and its cost, to be minimized, is determined by how much weight is shifted over which distances. On the other hand, a price function determines a price for some commodity at given points; nonexpansiveness of the price function means that no profit can be made from full-cost transport. The difference between the expected values of a price function under the given distributions is the profit to be made by having the commodity transported, and hence the amount one can offer to a logistics provider when outsourcing the transport.
@@
Both the coupling-based definition and the price-function-based definition have been generalized categorically to construct liftings of set functors to the category of (pseudo-)metric spaces [4] and quantitative lax extensions of set functors [39], where the latter are distinguished by applying to unrestricted quantitative relations instead of only to (pseudo-)metrics. Metric functor liftings and lax extensions in particular serve to give a general treatment of behavioural distances on quantitative systems such as probabilistic, weighted, or metric [2] transition systems in the framework of universal coalgebra [32]. In this framework, functors serve as parameters determining a type of systems as their coalgebras; for instance, coalgebras for the distribution functor are Markov chains. In the generalized setting, the coupling-based construction is often referred to as the Wasserstein lifting or extension, and the price-function-based one as the Kantorovich lifting or extension, but other names are found in the literature, such as coupling-based [23] and codensity [25] lifting, respectively. Both constructions are parametrized over a choice of quantitative modalities; the classical case involves, on both sides, only one modality, the expectation modality.
@@
The interest in having a price-function-based presentation of a given functor lifting or lax extension lies inter alia in the fact that one obtains a quantitative Hennessy-Milner property for the quantitative modal logic generated by the respective modalities [27, 39]. This property states coincidence of the behavioural distance induced by the given lifting or extension with the logical distance induced by the respective quantitative modal logic; thus, high distance between two states can always be certified by means of a modal formula, a prominent principle in the study of behavioural distances (e.g. [36, 10, 39, 27, 13] and more recently [31, 35]). It has been shown that every metric functor lifting that preserves isometries [14] and every quantitative lax extension [39] is Kantorovich, i.e. can be presented via the generalized price-function-based construction, using however a rather large (in particular typically infinite) set of modalities called the Moss modalities.
@@
An important open point that remains is thus the question of what we term generalized Kantorovich-Rubinstein duality: In which cases does a coupling-based distance given by a choice of modalities for a functor coincide with the price-function-based distance for the same modalities? Known positive examples include, as mentioned, the classical Kantorovich-Rubinstein distance on distributions, but also the Hausdorff distance on the powerset. We begin our analysis by giving an example of a natural coupling-based metric for which generalized Kantorovich-Rubinstein duality in this sense fails, namely the standard $p$-Wasserstein metric for $p>1$ (which minimizes $p$-th roots of the expectation of $p$-th powers of couplings). As our main contribution, we then provide two new positive examples, namely the Lévy-Prokhorov distance on probability distributions [30] and the standard distance on the convex powerset, whose elements are convex sets of probability distributions. The Lévy-Prokhorov distance has seen a recent rise in popularity due to its favourable robustness properties that make it suitable for tasks in machine learning (such as conformal prediction [3] and corruption resistance [5]); moreover, it has been shown in recent work [11] to induce precisely the behavioural distance defined by $\varepsilon$-bisimilarity [10]. In this case, the relevant modality is precisely the generally modality used in work on fuzzy description logics as an alternative formal correspondent of the natural-language term ‘probably’ [34]. The convex powerset plays a central role in the distribution semantics of Markov decision processes (or probabilistic automata) [7]. Its standard metric is just the composition of the Hausdorff and Kantorovich-Rubinstein metrics, and as such is again given via the coupling-based construction for a modality composed from the expectation modality and the standard fuzzy diamond modality [39]. We show that the generalized Kantorovich-Rubinstein duality holds w.r.t. this modality. Beyond the mentioned benefits regarding characteristic quantitative modal logics, we demonstrate in this case that the new price-function-based description also allows for more efficient computation of distances.
@@
Categorical distance constructions based on couplings were first considered in work on monoidal topology [21, 22]. The categorical price-function-based construction is predated by several constructions for specific functors in work on stochastic games [17]. Our price-function-based presentation of convex powerset complements earlier results on a coupling-based characterization [39] and a quantitative-algebraic presentation [28]. The modality underlying the distance on the convex powerset functor relates to Goubault-Larrecq’s previsions [16]; distances between such previsions are also expressed through a combination of (topological versions of) the Hausdorff and Kantorovich-Rubinstein-distances [17, 18, 19]. There has been recent work on what are termed correspondences between the generalized coupling-based and price-function-based constructions where a single modality is assumed for the former, while an associated set of modalities is considered for the latter [23]. Correspondences in this sense thus lie between dualities as considered here, where we insist on the same modalities being used on both sides, and general theorems on Kantorovich presentations of lax extensions [39] and classes of functors [14] that use very large sets of liftings. Results are obtained for functors constructed from the main known instances (distributions with the standard Kantorovich-Rubinstein metric, powerset) by applying coproduct and product. The price-function-based construction is sometimes referred to as the codensity construction [25, 26, 23], and as such has been used for logical characterizations of behavioural distances as mentioned above [26] but also for game characterizations [25]. The problem of generalized Kantorovich-Rubinstein duality has been stated already in work introducing coalgebraic behavioural distances [4], and a simple counterexample has been given; the counterexample we give here is distinguished by involving a quantitative modality that satisfies an analogue of two-valued separation [29].
@@ 2 Preliminaries
A function between the underlying sets $f\colon X\to Y$ of two pseudometric spaces $(X,d_{X})$, $(Y,d_{Y})$ is nonexpansive if distances are not increased by $f$, explicitly if for all $x,y\in X$ it holds that $d_{Y}(f(x),f(y))\leq d_{X}(x,y)$. Pseudometric spaces and nonexpansive functions between them form a category $\mathbf{PMet}$. The full subcategory of $\mathbf{PMet}$ spanned by metric spaces is denoted by $\mathbf{Met}$.
@@
Our main results derive some of their interest from their relevance to behavioural distances in coalgebras. Generally, the framework of universal coalgebra [32] is based on abstracting state-based systems as $F$-coalgebras for a functor $F\colon\mathbf{C}\to\mathbf{C}$ on a category $\mathbf{C}$, with $F$ determining the type of the system. Specifically, an $F$-coalgebra is a pair $(C,\gamma)$ consisting of a $\mathbf{C}$-object $C$, thought of as an object of states, and a morphism $\gamma\colon C\to FC$ determining transitions from states to structured collections of successor states, with the structure determined by $F$. A homomorphism between $F$-coalgebras $(C,\gamma)$ and $(D,\delta)$ is a $\mathbf{C}$-morphism $h\colon C\to D$ such that $\delta\circ h=Fh\circ\gamma$.
We list some common functors that will be useful in the further technical development and their associated coalgebras.
@@
    $\mathcal{D}X=\{\mu\colon X\to[0,1]\mid\mu(x)>0\text{ for finitely many }x\in X\text{ and }\sum_{x\in X}\mu(x)=1\}.$
@@
    The coalgebras of $\mathcal{D}$ are precisely (discrete-time) Markov chains.
@@ 3 Dual Characterizations of Metrics
    The Kantorovich-Rubinstein lifting $\overline{\mathcal{D}}\colon\mathbf{PMet}\to\mathbf{PMet}$ of $\mathcal{D}$ equips the set $\mathcal{D}X$ with the Kantorovich-Rubinstein distance $\delta^{\mathsf{KR}}(d_{X})$, which is defined as
    $\delta^{\mathsf{KR}}(d_{X})(\mu,\nu):=\sup\{\mathbb{E}_{\nu}({f})-\mathbb{E}_{\mu}({f})\mid f\colon(X,d_{X})\to([0,1],d_{e})\text{ nonexpansive}\},$
@@
    Let $\lambda$ be a predicate lifting for $F$. The price-function-based lifting of $F$ to the category of pseudometric spaces sends a metric $d_{X}$ to $K_{\lambda}(d_{X})\colon FX\times FX\to[0,1]$ defined by
@@
    Let $s\in FX$ and $t\in FY$. The set of couplings $\Gamma(s,t)$ is defined as
    $\Gamma(s,t):=\{c\in F(X\times Y)\mid F\pi_{1}(c)=s\text{ and }F\pi_{2}(c)=t\}.$
@@
    Let $\lambda$ be a well-behaved predicate lifting for $F$ and assume that $F$ preserves weak pullbacks. The coupling-based lifting of $F$ to the category of pseudometric spaces is defined as the lifting that equips $FX$ with $W_{\lambda}(d_{X})$ where
@@
While the price-function-based lifting assumes no conditions on supplied structures, the coupling-based lifting is significantly more particular, requiring both pullback preservation of the underlying functor and that the predicate lifting be well-behaved, to ensure that $W_{\lambda}(d_{X})$ is a pseudometric whenever $d_{X}$ is [4, 6].
@@
    Let $F=\mathcal{D}$ and let $\lambda$ be the predicate lifting calculating expected values: For $\mu\in\mathcal{D}X$ and $f\in[0,1]^{X}$ we have $\lambda_{X}(f)(\mu)=\mathbb{E}_{\mu}({f})$. Then $\delta^{\mathsf{KR}}=K_{\lambda}=W_{\lambda}$.
If, like in the two examples above, the categorical price-function-based and coupling-based constructions $K_{\lambda}$ and $W_{\lambda}$ coincide, we say that generalized Kantorovich-Rubinstein duality holds. This name is motivated by the particular case of Example˜3.2, the classical Kantorovich-Rubinstein duality that dates back to the beginnings of transportation theory [24]. It is important to note that generalized duality in this sense may fail, as demonstrated by the case of $p$-Wasserstein distance that we discuss later. It is a general fact [4, Theorem 5.27] that $K_{\lambda}\leq W_{\lambda}$, so duality hinges on the inequality $W_{\lambda}\leq K_{\lambda}$.
@@
The names of the two constructions are motivated by the probabilistic case as well. The coupling-based presentation is closely related to optimal transportation theory: In the case of the Kantorovich-Rubinstein lifting, one may view probability distributions $\mu,\nu$ as producers and consumers of a resource respectively (with the amount of the resource produced/consumed being fixed to $1$). Now one wants to transport the produced resources to the consumers in the economically most efficient possible way, minimizing the average distance each unit of resource needs to travel. Couplings $c$ of $\mu$ and $\nu$ can then be seen as transport plans, with $\mathbb{E}_{c}({d_{X}})$ giving the total cost of carrying out the plan. Then the Kantorovich-Rubinstein distance gives us by definition the minimal possible cost.
In the distribution case (Example˜3.2), the price-function-based presentation can be explained via the analogy of “outsourcing the transport” by defining a function $f$ that assigns a price to each $x\in X$. This function must satisfy the requirement that it is nonexpansive, i.e., that the difference of prices assigned to $x,y$ is always at most $d_{X}(x,y)$ (meaning that no extra profit can be made from such a transport). The overall profit under such a price function is then the income obtained from the consumers ($\mathbb{E}_{\nu}({f})$) minus the cost paid to the producers ($\mathbb{E}_{\mu}({f})$). Taking the maximum over all such $f$ gives us the value of the Kantorovich-Rubinstein lifting.
@@
The price-function-based lifting is closely related to characteristic multi-valued modal logics: These are logics in which formulae $\varphi$ receive semantics in coalgebras $(X,\gamma)$, inducing an interpretation function
$\llbracket\varphi\rrbracket_{\gamma}\colon X\to[0,1].$
The logical distance of two states is then the supremum of all distances witnessed by such formulas.
The semantics of a modal operator $L$ in these types of logics is usually given as a predicate lifting $\lambda\colon[0,1]^{-}\Rightarrow[0,1]^{F-}$, with $\llbracket L\varphi\rrbracket_{\gamma}$ being inductively defined as $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$. Then expressivity of the logic (the fact that the behavioural distance can be witnessed by formulae of the logic arbitrarily closely) can be shown by exploiting the price-function-based presentation of the lifting. In fact, the interpretation $\lambda_{X}(\llbracket\varphi\rrbracket_{\gamma})\circ\gamma$ can be viewed as emulating one step of the functional $\Phi_{\gamma}$, under the condition that the interpretations of formulae $\llbracket\varphi\rrbracket_{\gamma}$ are able to approximate any nonexpansive function $f\colon X\to[0,1]$ arbitrarily closely [13].
Hence, the duality can be used very fruitfully: One can use the coupling-based view to compute (an under-approximation of) the behavioural distance and switch to price functions to determine the distinguishing formula witnessing this distance.
@@
There also exists a parametrized version of the Kantorovich-Rubinstein distance, called the $p$-Wasserstein distance for some real parameter $p\geq 1$}. In categorical terms, it is given by the predicate lifting $\lambda_{p}(f)(\mu)=(\mathbb{E}_{\mu}({f^{p}}))^{\frac{1}{p}}$ (where $f^{p}$ takes the $p$-th power pointwise), that is:
@@
For $p=1$ this is just the usual Kantorovich-Rubinstein distance and duality holds. If $p>1$, however, then duality may fail, and the corresponding price-function-based construction $K_{\lambda_{p}}$ may be strictly below $W_{\lambda_{p}}$. This gives witness to the general idea from transportation theory that $p=1$ constitutes a special case among the family of Wasserstein distances:
@@
$W_{\lambda_{2}}(d)(\mu,\nu)=(\delta^{\mathsf{KR}}(d)(\mu,\nu))^{\frac{1}{2}}=(\tfrac{1}{3})^{\frac{1}{2}}=\tfrac{1}{\sqrt{3}},$
@@
$g(0)+g(1)=(\tfrac{1}{3}\cdot g(0)+\tfrac{2}{3}\cdot g(1))+(\tfrac{2}{3}\cdot g(0)+\tfrac{1}{3}\cdot g(1))\\ \leq g(\tfrac{1}{3}\cdot 0+\tfrac{2}{3}\cdot 1)+g(\tfrac{2}{3}\cdot 0+\tfrac{1}{3}\cdot 1)=g(\tfrac{1}{3})+g(\tfrac{2}{3}).$ (1)
Additionally, nonexpansiveness of $f$ implies that
$b^{2}-a^{2}=(b-a)(a+b)=(g(1)-g(0))(a+b)\leq(1-0)(a+b)=a+b.$ (2)
@@
[…]
Our earlier assumption that $a<b$ implies that $g(\tfrac{2}{3})+g(\tfrac{1}{3})$ is positive, so we can divide by it on both sides, which results in the claimed inequality. ∎